8. Find a generating function for the sequence (a0, a1,....ar.....), where ar=the number of non negative integral solutions to e1+e2 +....en=r where 0 less than or equal to e1 less than or equal to 1 for each i=1,2,3,.....n


1
Expert's answer
2022-03-07T04:57:01-0500

The number of non-negative integral solutions to e1+e2 +...+en=r, where "e_i\in\{0,1\}" is equal to the number of all r-subsets of the set "I_n=\{0,1,\dots,n\}".

The 1-to-1 correspondence can be given as following:

"A\subset I_n \leftrightarrow (e_1,e_2,\dots,e_n)\in\{0,1\}^n: e_i=1\, {\rm iff }\, i\in A"

The number of all r-subsets of the set "I_n" , is equal to the binomial coefficient "\binom{n}{r}".

Therefore, the generating function for the sequence "(a_0, a_1,\dots,a_r,\dots)=(\binom{n}{0},\binom{n}{1},\dots,\binom{n}{r},\dots)" is

"f(z)=\sum\limits_{r=0}^{\infty}a_rz^r=\sum\limits_{r=0}^{\infty}\binom{n}{r}z^r=(1+z)^n"

The final answer does not depend on "r", because, by definition, a generating function "f(z)" for the sequence (a0, a1,....ar.....) is the sum of the series "f(z)=\sum\limits_{r=0}^{\infty}a_rz^r", which does not depend on "r".

Answer. "f(z)=(1+z)^n"